OpenStudy (anonymous):

Consider the functions f(x) = 4x^2 and g(x) = 5(x + 7). Find f[g(−5)]"

OpenStudy (anonymous):

this is my new question. Last 1

OpenStudy (anonymous):

uh, this is pretty simple and ill solve it , but i want you to try it first

OpenStudy (anonymous):

Then ill post my solution

OpenStudy (anonymous):

i know how to do the first part but idk how to Find f[g(−5)]

OpenStudy (anonymous):

wait dont tell me the answer let me try first

OpenStudy (anonymous):

100[-60(-5)] then solve this?

OpenStudy (anonymous):

would the answer be 30000

OpenStudy (anonymous):

uh, nope... solve g(-5)

OpenStudy (anonymous):

Put (-5) into g(x), 5(x+7)

OpenStudy (anonymous):

400?

OpenStudy (anonymous):

Very Good

OpenStudy (anonymous):

Okay.. first step is to calculate the function in the bracket that is g(-5); = 5(-5+7) = 10 then sub the answer into the f(x); = 4(10)^2 = 400

OpenStudy (anonymous):

omg thanks!

OpenStudy (anonymous):

My pleasure

OpenStudy (anonymous):

yay i did it by myself!

OpenStudy (anonymous):

Lol yes you did

OpenStudy (anonymous):

sorta...

OpenStudy (anonymous):

you did well